OpenStudy (ksaimouli):

A=40 degree,a=17,b=11 slove triangle

OpenStudy (cwrw238):

use sine rule to find value of angle B: 17 / sin 40 = 11 / sin B

OpenStudy (ksaimouli):

i got all thei nees to verify r answer

OpenStudy (cwrw238):

having found angle B angle C = 180-(A+B)

OpenStudy (cwrw238):

sorry i dont understand what you mean - you want me to verify your answers?

OpenStudy (ksaimouli):

i got angle b 41 and angle c=90

OpenStudy (ksaimouli):

side c=45

OpenStudy (ksaimouli):

is these correct

OpenStudy (cwrw238):

ok sin B = 11 sin 40 / 17 = 0.4192 B = 24.8 degrees C = 180-(40+24.8) = 115.2

OpenStudy (ksaimouli):

wherethat 11 came from

OpenStudy (ksaimouli):

opps b=28

OpenStudy (ksaimouli):

thx

OpenStudy (cwrw238):

b = 11 in your q